OpenStudy (anonymous):

y+3=2/5(x+9) write this equation in standard form PLEASE HELP

OpenStudy (karatechopper):

standard form= Ax+By=C

OpenStudy (karatechopper):

first you want to distribute by using distributive property

OpenStudy (callisto):

Nah.. isn't it Ax+By+C=0? y+3=2/5(x+9) y+3 = (2/5)x + (2/5)(9) Both sides x5 5y +15 = 2x + 18 Swap the terms 2x+18 -(5y+15) =0 2x - 5y +3 =0

OpenStudy (anonymous):

It is actually Ax + By = C. Ax(^2) + By + C = 0 is more for quadratic equations.

OpenStudy (callisto):

Okay, thanks

Directrix (directrix):

Standard Form: the standard form of a line is in the form Ax + By = C where A is a positive integer, and B, and C are integers.

Directrix (directrix):

So, 2x - 5y = -3